# Rethinking Data Input for Point Cloud Upsampling

### v1

@@ 4.1 Discussion
More effective local feature extraction: When upsampling on AS method, each AS slice contains fewer point clouds, and the graph convolutional network can extract local features more finely. Due to each AS slice of the point cloud being within a small range, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD.
Set $\mathcal{P}_{k}$ represents the $k$-th AS slice, $\mathcal{P}_{k}=\{\mathbf{p}_{i}\mid i\in I_{k}\}$, where $I_{k}$ is the point index set for AS $\mathcal{P}_{k}$. For each AS slice $\mathcal{P}_{k}$, its point cloud can be represented as $\mathbf{X}_{k}\in\mathbb{R}^{N_{k}\times d}$, where $N_{k}$ is the number of AS slices and $d$ is the dimension of the points.
In a graph convolutional network, the output of the $l$-th layer is represented as:
$\mathbf{H}_{k}^{(l)}=\sigma(\mathbf{A}_{k}\mathbf{H}_{k}^{(l-1)}\mathbf{W}^{(l)})$
In an AS slice, the adjacency matrix $\mathbf{A}_{k}$ is denser, with tighter connections between points, making convolution operations more effective in capturing local features.
Local consistency of AS slices: When dealing with smaller AS slices, the local consistency of point clouds is stronger, that is, the similarity between points is higher. This enables graph convolutional networks to better utilize these local consistency for learning and reconstruction, reducing local errors introduced by Patch based method.
The AS method deals with point clouds with strong local consistency, which can be measured by the local covariance matrix.
For AS slice $\mathcal{P}_{k}$, its point cloud $\mathbf{X}_{k}$ have the local covariance matrix is:
$\mathbf{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}(\mathbf{x}_{i}-\mathbf{\mu}_{k})(\mathbf{x}_{i}-\mathbf{\mu}_{k})^{\top}$
Where $\mathbf{\mu}_{k}$ is mean vector of midpoint of AS slice $\mathcal{P}_{k}$ :
$\mathbf{\mu}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}$
Local covariance matrix $\mathbf{C}_{k}$ can measure the local consistency of point clouds. In the case of strong local consistency, the smaller feature values of $\mathbf{C}_{k}$ make local feature extraction more effective.
@@
4.2 Future Work
In point cloud 3D model sampling, such as Pu-Edgeformr++ [10], it can achieve good point cloud edge recognition and sampling. However, taking the PU-GCN model used in this article as an example, the existing point cloud upsampling models are not convincing in their performance on the ABC dataset. The method of reverse engineering CAD models follows a series of typical steps: data capture and preprocessing, point cloud segmentation, face classification, and analysis of faces generated by segmentation methods [3]. The poor quantitative and qualitative performance on the CAD dataset ABC may be some challenges due to the heterogeneous and unstructured nature of manufacturing data [16]. In the future, the application of point cloud upsampling in reverse engineering and CAD datasets may require a similar approach to Point2CAD [14], which involves dividing point clouds into clusters corresponding to CAD model faces, fitting each cluster with an analytical surface primitive, and proposing a novel fitting scheme for free-form surfaces based on neural representation.
5 Conclusion
In order to ensure the model integrity of point cloud upsampling, this article mainly studies the input of point cloud data. Specifically, this article proposes a new data input method, also known as a new training model method, called Average Segment. And experiments were conducted on a stable and lightweight PU-GCN model. The traditional patch method still performs slightly better in both CD and HD indicators, but through relevant ablation research and visualization, we can see that the AS method can maintain local features within the model. In future work, we consider combining point cloud models with mesh face clusters to perform pairwise surface intersections to identify object edges, thereby achieving better edge, local, and global representations.

### v2

@@ 4. Discussion
5. Future Works
PU-GCN model used in this article as an example, the existing Patch based upsampling method is better than AS method we proposed. This may be due to the difference in dataset size between the two methods. In future work, we will delve deeper into the differences in interpretability between Patch and AS methods through representation learning methods. At the same time, due to the similarity between the use of point clouds and the completion of point clouds. We will refer to the method in article  (Xia et al., 2021) and use encoders for the data input of Patch and AS respectively, in order to merge global and edge features and achieve better results.
6. Conclusion
To ensure the model integrity of point cloud upsampling, this article mainly studies the input of point cloud data. Specifically, this article proposes a new data input method, also known as a new training model method, called Average Segment. And experiments were conducted on a stable and lightweight PU-GCN model. The traditional patch method still performs slightly better in both CD and HD indicators, but through relevant ablation research and visualization, we can see that the AS method can maintain local features within the model.